Two vehicles are approaching an intersection. One is a 2300 kg pickup traveling at 18.0 m/s from east to west (the −x- direction), and the other is a 1600 kg sedan going from south to north (the +y− direction at 24.0 m/s ).

What is the direction of the net momentum? (  ∘ west of north.)
Answer #1

The x-component is all in the pickup:

px = 2300kg * -18.0m/s = -41 400 kg·m/s

py = 1600kg * 24.0m/s = 38 400 kg·m/s

(a) |p| = √(px² + py²) = 56,467 kg·m/s

(b) Θ = arctan(py/px) = arctan(38,400/-41,400) = -42.85º, or 42.85º N of West,

= 90 – 42.85 = 47.15 degrees west of north

Kindly rate :)